OpenStudy (anonymous):

8 1/6 divided by 4 2/3

OpenStudy (anonymous):

Convert the two mixed fractions into improper fractions first

OpenStudy (anonymous):

1.75 or 7/4

OpenStudy (anonymous):

\[\frac{ 49 }{ 6 } \div \frac{ 14 }{ 3 }\]

OpenStudy (anonymous):

then flip the second fraction and change the sign to multiply

OpenStudy (anonymous):

\[8 + \frac{ 1 }{ 6 } \div 4+ \frac{ 2 }{ 3 }\] That's what your equation looks like so far ... Convert your 2 mixed fractions into improper fractions then flip it over and multiply. Your answer should look like this : \[\frac{ 7 }{ 4}\] Hope this helped!